Find the exact value for each trigonometric function.
cos 675


8. csc (-240)


9. cot 900



10. tan 570


11. sec (-780)


12. csc 90



Answer by lwsshak3(11628)

Find the exact value for each trigonometric function.
cos 675
675-360=315º=reference angle of 45º in quadrant IV where cos>0
cos 675º=cos 45º=√2/2
8. csc (-240)
240-180=60º reference angle in quadrant II where csc <0
csc(-240)=csc 60º=-2/√3=-2√3/3
9. cot 900
900-720=180º
cot 900=cot 180=0
10. tan 570
570-360=210=reference angle of 30º in quadrant III where tan >0
tan 570=tan 30º=√3/3
11. sec (-780)
780-720=60º reference angle in quadrant IV where sec<0
sec (-780)=sec 60º=-2
12. csc 90=1/sin 90º=1/1=1
Note: When given angle is negative rotate clockwise, when positive, rotate counter clockwise.